OpenStudy (bloomlocke367):

help! How would I graph this using a graphing calculator? \( 7^x>3^x\)

OpenStudy (bloomlocke367):

@IrishBoy123

OpenStudy (solomonzelman):

this equation just contains the x, and therefore I am really suspicious that this is not the initial question

OpenStudy (solomonzelman):

Or is it still the initial question - to graph \(7^x>3^x\) ?

OpenStudy (bloomlocke367):

it is tho.. I can pull up the review packet online and screenshot it.

OpenStudy (bloomlocke367):

http://prntscr.com/9mkb6k

OpenStudy (solomonzelman):

oh that is a little different

OpenStudy (solomonzelman):

to "solve the system graphically.

OpenStudy (solomonzelman):

so you are, I think, graphing two functions \(\color{#000000 }{ \displaystyle f(x)=7^x }\) \(\color{#000000 }{ \displaystyle g(x)=3^x }\) (I am labelling them f(x) and g(x) for convinience) And you want to know for which values of x is \(\color{#000000 }{ \displaystyle 7^x>3^x }\)

OpenStudy (solomonzelman):

Well, I mean to say: you want to know for which values of x is \(\color{#000000 }{ \displaystyle f(x)>g(x) }\), but this is really the same thing

OpenStudy (solomonzelman):

((Although there is a much easier way to tell without graphing))

OpenStudy (bloomlocke367):

so.. I graph them separately?

OpenStudy (bloomlocke367):

@SolomonZelman, sorry I had to do something for my dad.

OpenStudy (solomonzelman):

yes, graph them separately, and tell me where f(x)>g(x)

OpenStudy (bloomlocke367):

oki. One second

OpenStudy (bloomlocke367):

I'm not sure what I'm looking for tbh .-.

OpenStudy (bloomlocke367):

wait. I see now XD

OpenStudy (bloomlocke367):

when x is greater than 0. I had to zoom in on my calculator

OpenStudy (bloomlocke367):

Thanks for helping :D

OpenStudy (solomonzelman):

YW

OpenStudy (anonymous):

\[7^x>3^x~only~if~x>0\]

OpenStudy (solomonzelman):

You can also tell this not graphically: let \(c\) be any positive number (even really really close to 0): \(\color{#000000 }{ \displaystyle 7^c>3^c }\) \(\color{#000000 }{ \displaystyle 7^0=3^0 }\) \(\color{#000000 }{ \displaystyle 7^{-c}~~~~~?~~~~~3^{-c} }\) \(\color{#000000 }{ \displaystyle \frac{1}{7^c}~~~~~?~~~~~\frac{1}{3^c} }\) \(\color{#000000 }{ \displaystyle \frac{1}{7^c}~~~~~<~~~~~\frac{1}{3^c} }\) (you are dividing by larger value on the right, therefore the result is smaller)

OpenStudy (solomonzelman):

by -c, I am denoting a negative number. (Since c is a positive number)

OpenStudy (solomonzelman):

Oh m,y bad, you are dividing by a larger value on the left (not right)

OpenStudy (solomonzelman):

confused my left and rght :D

OpenStudy (bloomlocke367):

haha, it's okay. thanks!

OpenStudy (solomonzelman):

yw